Find the optimal solution for the following problem. Maximize C = 4x + 12y + 3z subject to 15x + 5y + 12z ≤ 75 6x + 2y + 8z ≤ 150 and x ≥ 0, y ≥ 0.
Answer #1

Given that,

Maximize C = 4x + 12y + 3z subject to 15x + 5y + 12z ≤ 75 6x + 2y + 8z ≤ 150 and x ≥ 0, y ≥ 0 and z ≥ 0,

Here, x = X1, y = X2, z = X3, C = Z,

Calculation:

Pivot row (Row 1):
75 / 5 = 15
15 / 5 = 3
5 / 5 = 1
12 / 5 = 12 / 5
1 / 5 = 1 / 5
0 / 5 = 0

Row 2:
150 - (2 * 15) = 120
6 - (2 * 3) = 0
2 - (2 * 1) = 0
8 - (2 * 12 / 5) = 16 / 5
0 - (2 * 1 / 5) = -2 / 5
1 - (2 * 0) = 1

Row Z:
0 - (-12 * 15) = 180
-4 - (-12 * 3) = 32
-12 - (-12 * 1) = 0
-3 - (-12 * 12 / 5) = 129 / 5
0 - (-12 * 1 / 5) = 12 / 5
0 - (-12 * 0) = 0

or, as per question, x= 0, y=15, z=0, Maximize C = 180